OpenStudy (anonymous):

how do i decompose a function using trig of an arc trig function. Ex. Square root of 1-x^2

OpenStudy (loser66):

don't understand

OpenStudy (anonymous):

my text book says "decompose each of the following algebraic functions by using it as a trig function of an arctrig function

OpenStudy (anonymous):

\[\sqrt{1- {x}^{1}}\]

OpenStudy (alekos):

standby, i'm working on it

OpenStudy (loser66):

thanks @alekos

OpenStudy (anonymous):

thanks! and if you can explain it, that would help so much!

OpenStudy (alekos):

this is quite tricky but here it goes... y = sqrt(1 - x^2) we let x = sin(theta) i.e. theta = arcsinx so we have y = sqrt[1 - sin^2(theta)] follow so far?

OpenStudy (anonymous):

yes

OpenStudy (alekos):

OK, from this we can see that y = cos(theta) due to the relationship sin^2(theta) + cos^2(theta) = 1 so we have cos(theta) = sqrt(1-x^2) but theta = arcsinx from above and hence we have cos(arcsinx) = sqrt(1-x^2)

OpenStudy (anonymous):

holy crap... i never saw it like that

OpenStudy (alekos):

we can also deduce that sin(arccosx) = sqrt(1-x^2)

OpenStudy (alekos):

yeah, that's how its done. quite tricky and hard to get your head around

OpenStudy (anonymous):

yea, my book, nor my teacher never helps... thanks!

OpenStudy (alekos):

thats ok, your welcome